Both solution X and solution Y contain alcohol and water. In solution X the ratio
of alcohol to water is 3:2. When equal amounts of solution X and solution Y are
mixed, the ratio of alcohol to water is 3:4. What is the ratio of alcohol to water in solution Y? Write your answer in the form A:W.

Found 2 solutions by richwmiller, vksarvepalli:
Answer by richwmiller(17219)

I'll try this one because it is a little different from the other mixture problems.
solution x ratio is 3:2 making it 3 /5 alcohol
solution y is 3:4 making it 3/7 alcohol
Let the amount of solution x and solution y be a
(3/5)a+ya=(3/7)*2a
a21/35 +35ay/35=6*5a/35
multiply by 35
21a+35ay=30a
divide by a
21+35y=30
35y=9
y=9/35
so the solution y is 9/35 alcohol and 26/35 water
so the ratio is 9:26 A:W



Answer by vksarvepalli(154)

let the ratio of alcohol to water in Y be a:b
now suppose 5 litres each of both the sol's are mixed (you can take any value it doesn't matter just they have to be equal in volume)
then we will have two equations in a and b
a+b=5
and 3b-4a=6
solving these we get the answer a:b=9:26